# Early Risk Prediction of Pediatric Cardiac Arrest from Electronic Health Records via Multimodal Fused Transformer

### v2

@@ I Introduction
Cardiac arrest (CA) in the pediatric population is a critical and life-threatening event with significant implications for morbidity and mortality [1]. While less common than in adults, pediatric CA is often associated with a preceding period of physiological instability, making timely identification and intervention crucial. Many cases of CA in children occur in intensive care units (ICUs) [2], particularly cardiac ICUs [3], where children with congenital or acquired heart diseases are managed. In these high-stakes environments, the interplay of complex cardiac conditions, surgical interventions, and critical care therapies, coupled with significant cognitive and data overload, creates a unique set of challenges for clinicians [4]. Early prediction of CA [5] in this setting could enable proactive measures to prevent arrests and provide situational awareness. This underscores the importance of developing robust predictive models tailored to the pediatric cardiac ICU population, integrating clinical, physiological, and potentially novel data sources.
@@ II-A Problem Definition
We aim to employ a predictive model $f_{\theta}(\cdot)$ using a patient EHR data $\mathbf{x}_{t\leq\tau}$ available in the first $\tau$ hours of admission, to predict the probability of developing cardiac arrest $\hat{y}=f_{\theta}(\mathbf{x}_{t\leq\tau})$, where $\hat{y}\in\{0,1\}$ during the rest of the admission stay.
@@ II-B Data Source and Cohort Selection
CHOA-CICU database is a private pediatric intensive care database extracted from electronic health record for all pediatric patients (age less than 18 years) admitted from 1/1/2018 to 12/31/2023 to a large, quaternary, academic Pediatric Cardiac Intensive Care Unit (CICU) at Children’s Hospital at Atlanta and Emory University. We obtain 9 static risk factors from CHOA EHR, including: “gender, ethnicity, first race, second race, primary language, respiration distress, age, admission diagnoses, admission ICD 19 codes”. Among 184 temporal risk factors, we obtain 129 vital signs and nursing assessments from flow sheet after removing non-relevant items and removing missing rate $>0.5$ items. We also obtain 49 lab test results after removing missing rate $>0.5$ items, and 5 medications after removing missing rate $>0.7$ items. Table I briefly presents the number of patients included in the two data sources.
@@ IV-D Feature Importance Analysis
We use permutation importance method [37] to calculate the feature importance scores for our proposed perturbed copy of the data where this feature’s values have been randomly shuffled across rows. We set number of different permutation shuffles as $5$, and the confidence level as $0.95$. Top 15 important features with $p\_value\leq 0.05$ are presented in Fig. 3. Our analysis reveals that key clinical indicators, such as capillary refill, total bilirubin level, and hematocrit, are among the most important predictors for pediatric cardiac arrest. Notably, factors related to hemodynamic stability and oxygenation/ventilation—including CO2 level on arterial blood gas, and FIO2,—also rank highly, emphasizing their critical role in patient monitoring. Furthermore, the inclusion of demographic factors such as sex and race highlights the multifactorial nature of cardiac arrest risk in the pediatric population. Overall, this feature importance analysis not only identifies clinical risk factors but also uncovers novel predictors that may enhance early risk detection and guide future targeted interventions.

### v3

@@ I Introduction
[…]
@@ II-A Problem Definition
[…]
@@ II-B Data Source and Cohort Selection
CHOA-CICU database is a private pediatric intensive care database extracted from electronic health record for all pediatric patients (age less than 18 years) admitted from 1/1/2018 to 12/31/2023 to a large, quaternary, academic Pediatric Cardiac Intensive Care Unit (CICU) at Children’s Hospital at Atlanta and Emory University. We obtain 9 static risk factors from CHOA EHR, including: “sex, ethnicity, first race, second race, primary language, respiration distress, age, admission diagnoses, admission ICD-9 codes”. Among 184 temporal risk factors, we obtain 129 vital signs and nursing assessments from flow sheet after removing non-relevant items and removing missing rate $>0.5$ items. We also obtain 49 lab test results after removing missing rate $>0.5$ items, and 5 medications after removing missing rate $>0.7$ items. Table I briefly presents the number of patients included in the two data sources.
@@ IV-D Feature Importance Analysis
We use permutation importance method [38] to calculate the feature importance scores for our proposed perturbed copy of the data where this feature’s values have been randomly shuffled across rows. We set number of different permutation shuffles as $5$, and the confidence level as $0.95$. Top 15 important features with $p\_value\leq 0.05$ are presented in Fig. 3. Our analysis reveals that key clinical indicators, such as capillary refill (CRT), total bilirubin level, and hematocrit, are among the most important predictors for pediatric cardiac arrest. Notably, factors related to hemodynamic stability and oxygenation/ventilation—including CO2 level on arterial blood gas, and FIO2,—also rank highly, emphasizing their critical role in patient monitoring. Furthermore, the inclusion of demographic factors such as sex highlights the multifactorial nature of cardiac arrest risk in the pediatric population. Overall, this feature importance analysis not only identifies clinical risk factors but also uncovers novel predictors that may enhance early risk detection and guide future targeted interventions.